OpenStudy (anonymous):

help ASAP if a student predicted they would spend 125$ shopping but they actually spent $87.50 what is the percent of error? 42.9% 37.5% 30% 3%

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

ok dont hate me if im wrong but i think its 37.5

OpenStudy (anonymous):

The difference is 125 - 87.50 = 37.50 That is the amount of money they have to pay more. \(\dfrac{37.5*100\%}{125}= 30 \%\)

OpenStudy (anonymous):

well i was close :P

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

thank you